completeness and clarity

2. Find a particular solution for the IVP y - 2y + 2y = 8(t - 5), y(0) = 0, y0) = 1
Answer #1

Here we proceed through laplace transform.

If laplace transform of y +Y(s) then £ {y}+sY(s) – sy(0) - y0) and for {y} +sY(s) – y(0) Here take laplace transform of dif

-58 For £-1 52 - 2s + 2) we know £-1{eºsF(s)} = H(t - a)f(t – a) If £-{F(s)} = f(t) Using this formula £-1 2s + 2 b S2 – 2s +